# Symmetry breaking in two-dimensional turbulence induced by out-of-equilibrium fluxes

### v1

@@ Abstract
Considering the above boundaries in the ”thermodynamic” infinite domain limit of $L\gg l_{f}$, while keeping $(l_{\nu}/l_{f})^{2}\gg l_{f}/L$, only two phases remain: A-Jets for $l_{\nu}/l_{f}\gtrsim 0.4$ (coexistence of four types: horizontal/vertical jets, left/right bent jet) and a coexistence phase between A-Jets and a dipole for $l_{\nu}/l_{f}\lesssim 0.4$. For $l_{\nu}/l_{f}\gtrsim 0.4$ where only jets exist, any particular state breaks the $\pi/2$ symmetry of the domain. However, the symmetry could be statistically restored if the system randomly switches between states. We have not observed such transitions for A-Jets. Instead, we suggest that the systems symmetry is spontaneously broken. Indeed, a transition from e.g. horizontal jets to vertical ones requires closing the condensate voids and reopening them in the other direction. However, in the A-Jet state the length of the condensate voids parallel to the jets is extensive ($\propto L$), while the size of typical fluctuations is $l_{f}$. Therefore, closing the voids requires a fluctuation whose probability is expected to be exponentially suppressed as $l_{f}/L\to 0$. This ends our story of the LQG condensate: we found it is shaped by out-of-equilibrium fluxes, and that when the inverse inertial range is much longer than the direct one, the symmetry is spontaneously broken in the thermodynamic limit.

### v2

@@ Abstract
Then, for $l_{\nu}/l_{f}\gtrsim 0.4$ where only jets exist, any particular state breaks the $\pi/2$ symmetry of the domain. This symmetry could be statistically restored if the system randomly switches between states, but we have not observed such transitions for A-Jets. Instead, we suggest that the systems’ symmetry is spontaneously broken. Indeed, a transition from e.g. horizontal jets to vertical ones requires closing the condensate voids and reopening them in the other direction. However, in the A-Jet state the length of the condensate voids parallel to the jets is extensive ($\propto L$), while the size of typical fluctuations is $l_{f}$. Therefore, closing the voids requires a fluctuation whose probability is expected to be exponentially suppressed as $l_{f}/L\to 0$, implying spontaneous symmetry breaking in the thermodynamic limit.
Conclusion We found an unexpectedly rich landscape of large-scale flow configurations, absent in 2D Navier-Stokes. Its existence can be traced to the enslavement of the small-scale (direct) flux to the large-scale (inverse) flux occurring in LQG Svirsky et al. (2023a). We expect similar effects for flows in the regime $\rm{Fr}^{2}\ll\rm{Ro}\ll\rm{Fr}\ll 1$ when $l_{f}>L_{d}$, as the dynamics on scales larger than $L_{d}$ then follow LQG Polvani et al. (1994); Smith et al. (2002). The range of interactions is thus a relevant control parameter shaping coherent structures; when it is small enough, the interplay between out-of-equilibrium fluxes imposes a novel organizational principle.
Acknowledgements.